Sally drove from Town A to Town B at an average speed of 20 miles per hour. She then returned to Town A along the same road at an average speed of 30 miles per hour. What was Sally’s average speed, in miles per hour, for the entire trip?

Found 2 solutions by ewatrrr, kintosh88:
Answer by ewatrrr(24785)
Answer by kintosh88(1)

Ok so with this you need to do a little algebra to get you answer.
we are missing the time and distance.
let just say the distance for both trip is d+d or 2d
Now using the rate problem we will get the time by itself
d = r.t and t= d/r
so now we will find time for both trips : t = d/r + d/r
t= d/20 +d/30
= 3d/60 + 2d/30
= 5d/60
= d/12

Now that we have the time lets find the average distance
Average distance = distance / time
= 2d / d/12
= 2d x12/d
=24d/ d
your answer will be 24.